This problem has me perplexed. Help me please!
How many pounds of a chicken feed that sells for $5 per pound must be mixed with
400 pounds of a feed that sells for $8 per pound to make a chicken feed that sells for $7.50?

Answer by checkley71(8403)

5X+400*8=7.5(400+X)
5X+3200=3000+7.5X
5X-7.5X=3000-3200
-2.5X=-200
X=-200/-2.5
X=80 POUNDS OF CHICKEN IS NEEDED.
PROOF.
5*80+3200=7.5(480)
400+3200=3600
3600=3600